$SPAD/src/input mult3d.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

\getchunk{license}
)clear all
s := create3Space()$ThreeSpace(DFLOAT)
draw(curve(sin(t),cos(t),0),t=0..2*%pi,tubeRadius==.3, space == s)
draw(curve(sin(t)*cos(%pi/2),cos(t),0-sin(t)*sin(%pi/2)),_
     t=0..2*%pi,tubeRadius==.3, space == s)
v1 := draw(curve(sin(t),cos(t),0),t=0..2*%pi,tubeRadius==.3)
s1 := subspace(v1)
v2 := draw(curve(sin(t)*cos(%pi/2),cos(t),0-sin(t)*sin(%pi/2)),_
           t=0..2*%pi,tubeRadius==.3, space == s1)
s2 := subspace(v2)
)clear all
v := draw(surface(x*cos(y),x*sin(y),y*cos(x)),x=-4..4,y=0..2*%pi)
s := subspace(v)
makeObject(curve(4*sin(t),4*cos(t),-4),t=0..2*%pi,tubeRadius==.3, space == s)
makeObject(curve(4*sin(t),4*cos(t),0),t=0..2*%pi,tubeRadius==.3, space == s)
draw(curve(4*sin(t),4*cos(t),4),t=0..2*%pi,tubeRadius==.3, space == s)
)clear all
s := makeObject(surface(cos(t)*sin(p),sin(t)*sin(p),cos(p)),_
                t=0..2*%pi,p=0..%pi)
makeObject(curve(4*sin(t)*cos(%pi/2),4*cos(t),0-sin(t)*sin(%pi/2)),_
           t=0..2*%pi,tubeRadius==.3, space == s)
draw(curve(2*sin(t),2*cos(t),0),t=0..2*%pi,tubeRadius==.3, space == s)
)clear all
s := makeObject(curve(4*sin(t),4*cos(t),0),t=0..2*%pi,tubeRadius==.3)
makeObject(surface(cos(t)*sin(p),sin(t)*sin(p),cos(p)),_
           t=0..2*%pi,p=0..%pi, space == s)
makeObject(surface(cos(t)*sin(p)-3,sin(t)*sin(p)-3,cos(p)+2),_
           t=0..2*%pi,p=0..%pi, space == s)
draw(surface(cos(t)*sin(p)-6,sin(t)*sin(p)-6,cos(p)+4),_
     t=0..2*%pi,p=0..%pi, space == s)
)clear all
s := create3Space()$ThreeSpace(DFLOAT)
draw(cos(x*y)+1,x=-3..3,y=-3..3, space == s)

draw(-cos(x*y)-1,x=-3..3,y=-3..3, space == s)
)clear all
torusInKnot(p,q,r) ==
free sp
sp := create3Space()$ThreeSpace(DFLOAT)
a := 1
b := 1
ab := a+b
fac := 2*ab/(1.1*ab-sin(q*t))
Title := concat("Torus Knot of Type (", concat(p::String,
                concat(",", concat(q::String, ")"))))
Curve := curve(a*fac*cos(p*t), a*fac*sin(p*t), b*fac*cos(q*t))
r = 0 => draw(Curve, t=0..2*%pi, title == Title)
draw(Curve, t=0..2*%pi, tubeRadius == r, title == Title, space == sp)
torusInKnot(15,17,.1)
draw(curve(2.2*sin(t),2.2*cos(t),0),t=0..2*%pi,_
   tubePoints==10,tubeRadius==.7, space == sp)

———-

References^{[1] nothing}

5